A certain system can experience three different types of defects. Let A (i1, 2, 3) denote the event that the system has a defect oftype i. Suppose that: P(A)-0.12; P(A2)-0.07; P(A3)-0.05 a. b. c. d. What is the probability that they system does not have a type 1 defect? What is the probability that the system has both type 1 and type 2 defects? What is the probability that the system has both type 1 and type 2 defects but not a type 3 defect? What is the probability that the system has at most two of these defects?
Answer #1

a)

P(not have type I defect)=1-P(A)=1-0.12=0.88

b)P(A 1 n A2)=P(A1)+P(A2)-P(A1 n A2)=0.12+0.07-0.13=0.06

c)P(A1 n A2 n A3')=P(A1 n A2)-P(A1 n A2 n A3) =0.06-0.01=0.05

d)P(at most two defects)=1-P(all three defects)=1-0.01 =0.99